OpenStudy (anonymous):

How much will you save if you buy an item listed at $175.50 at a 35 percent discount?

OpenStudy (anonymous):

35% of 175.50

OpenStudy (anonymous):

$61.43

OpenStudy (anonymous):

35% = 35/100 "of" means to multiply, so \[\large \frac{35}{100} \times 175.50\]